OpenStudy (anonymous):

Solve the linear equation using substitution x=17-4y y=x-2 (Can you show steps please?)

OpenStudy (anonymous):

Solve the linear system*

OpenStudy (anonymous):

Well to solve a system of linear equations, we first need to write everything in terms of one variable. We shall write it in terms of (y) since they kindly isolated (x) for us already: x=17 - 4y we can just plug that in for (x) in the next equation: y=x-2 y = (17 - 4y) -2 y = 17 - 4y - 2 y + 4y = 15 5y = 15 y = 3

OpenStudy (anonymous):

Now we can take the y=3 and use it to find (x) x = 17 - 4y x = 17 - 4(3) x = 17 - 12 x = 5

OpenStudy (anonymous):

Thank you . :)